Circular No. 6035

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) R. Schulz, T. Encrenaz, J. A. Stuewe and G. Wiedemann communi-

cate: “We report preliminary results from our study of the impact sites of fragments A, C, D, E, F, G and H. Near-infrared spectra were recorded with the IRSPEC spectrometer at the European Southern Observatory’s 3.5-m NTT with a 4′′.4 slit aligned along the parallel through the impact sites. On July 17.04 UT the 2.12-μm H_{2} quadrupole line was detected with a resolving power of 2000 at the position of fragment A (∼ 5 hr after impact) as it moved along the jovian disk. On July 17.94 H_{3}^{+} was detected in four transitions around the strong 3.534-μm line with a resolving power of 1700. The region of the emission appears significantly extended in longitude with respect to the site of impact C (∼ 15.5 hr after impact). We continued to monitor the H_{3}^{+} multiplet at 3.5 μm during July 18.00–18.10 and 18.83–18.96. No evidence for H_{3}^{+} emission was found near the E–F impact region, whereas it was detected at the position of fragments D (∼ 14 hr after impact) and G (∼ 12 hr after impact). A weak H_{3}^{+} emission is visible on the leading side of impact region H (∼ 15 min after impact), but no H_{3}^{+} was detected in the region itself. On July 18.107 we started to monitor the E–F impact region between 2.02 and 2.05 μm. These observations allow the E and F sites to be distinguished. The spectrum of impact F (∼ 2 hr after impact) is characterized by a featureless continuum (which, if thermal, indicates temperatures of ∼ 1000 K), whereas the already evolved site E (∼ 11.5 hr after impact) shows little continuum but strong H_{2} S(1) line emission at 2.12 μm. These measurements suggest that H_{3}^{+} and H_{2} are present at more evolved impact sites rather than shortly after an impact.”

N. V. G. Sarma and the observing team at the 10.4-m millimeterwave telescope of the Raman Research Institute, Bangalore, report: “We detected short-duration increases in the jovian continuum flux density at 86 GHz following several of the impacts. Impact E: 400-percent increase on July 17.637 UT, the upper limit for the duration being 3 min as observations were made in scanning mode. Impact K: 50-percent increase on July 19.436 for 40 s, followed immediately by a 175-percent increase for ∼ 10 s. Impact N: increases by 87 percent and 125 percent on July 20.428 and 20.429, respectively, each lasting ∼ 20 s. Impact P: possible increase by 100 percent on July 20.629 for ∼ 3 s. Impact S: 70-percent increase on July 21.633 and 60-percent increase on 21.640, each for 30 s. As all the observations were made under monsoon weather conditions, some of them may be uncertain. We welcome confirmation from other millimeter-wave observatories.”

1994 July 22 Brian G. Marsden